The tiara worn by Kate Middleton for her wedding to Prince William of England contains 888 diamonds and belongs to the British monarchy. If each diamond in the tiara is 1.0 carat, and given that diamond is a form of carbon and that 1 carat is defined as 0.200 g, calculate the number of atoms in the gemstones of that tiara.
Answer #1

1 carat = 0.2 g, 888 carat= 888*0.2 = 177.6 g
Molar mass of carbon= 12 g/mol
hence 12 g of Carbon contains 1 mole of carbon atoms
so 177.6 g of carbon contains 177.6/12 = 14.8 mole of carbon atoms = 14.8*6.023x1023 carbon atoms = 8.91404x1024 carbon atms.
this is number of atoms in the gemstones of that tiara.